One of the largest barometers ever built was an oil-filled barometer constructed in Leicester, England in 1991. The oil had a height of 12.2 m. Assuming a pressure of 1.013 × 105 Pa, what was the density of the oil used in the barometer? Test Bank, Question 23 One of the largest barometers ever built was an oil-filled barometer constructed in Leicester, Englanduming a pressure of 1.013 × 105 Pa, what was the density of the oil used in the barometer?
Answer #1

Given data,

pressure = 1.013 x 105

Height = 12.2 m

We know that,

Pressure = \rho g h

1.013 * 105 = density x 9.81 x 12.2

Density = 1.013 * 105 / 9.81 x 12.2

= 101300 / 119.682

= 846.40 kg / m3

Option 3 is correct